### v1

# SMARTe-VR: Student Monitoring and Adaptive Response Technology for e-learning in Virtual Reality

Abstract
This work introduces SMARTe-VR, a platform for student monitoring in an immersive virtual reality environment designed for online education. SMARTe-VR is aimed to gather data for adaptive learning, focusing on facial biometrics and learning metadata. The platform allows instructors to create tailored learning sessions with video lectures, featuring an interface with an Auto QA system to evaluate understanding, interaction tools (e.g., textbook highlighting and lecture tagging), and real-time feedback. Additionally, we release a dataset containing 5 research challenges with data from 10 users in VR-based TOEIC sessions. This dataset, spanning over 25 hours, includes facial features, learning metadata, 450 responses, question difficulty levels, concept tags, and understanding labels. Alongside the database, we present preliminary experiments using Item Response Theory models, adapted for understanding detection using facial features. Two architectures were explored: a Temporal Convolutional Network for local features and a Multilayer Perceptron for global features.
1 Introduction
In recent years, education has undergone a transformation driven by the digital revolution. A key factor in this change has been the growing adoption of online education or e-learning, further accelerated by the COVID-19 pandemic, which forced a transition toward remote teaching models. This modality is expected to continue growing exponentially over the next 20 years (Chen 2018; Daza et al. 2023b).
Online education has traditionally been delivered through devices such as PC, laptops, and even tablets. However, as emerging technologies like virtual reality (VR) have become more affordable in recent years, a question arises: could these environments be more engaging than traditional e-learning? Through the metaverse, VR offers a sense of presence and interactivity that has often been missing in online education (Nagao 2023).
One of VR’s key strengths is its ability to more easily capture biometric and behavioral data, thanks to sensors placed on the face, hands, and other areas. This level of data collection is challenging to achieve in conventional e-learning environments. However, VR also presents challenges like fatigue and cybersickness after prolonged use, raising questions about its viability in education (Souchet et al. 2023).
This data collection capability not only facilitates the development and adoption of new learning systems based on biometric signal analysis, behavioral data, and learning analytics, but also addresses specific challenges in online education. Among these challenges is the lack of direct contact between teachers and students, which complicates aspects such as personalized learning (Kebritchi, Lipschuetz, and Santiague 2017).
VR has the potential to enhance educational quality through the implementation of Adaptive Learning (AL) systems (Nagao 2019; Barbosa et al. 2024). Learning analytics tools, such as Item Response Theory (IRT), play a crucial role in these systems, as they enable continuous and personalized assessment of students’ knowledge and performance (Zhou, Suzuki, and Kumano 2024).
The emotional and cognitive state of a student directly impacts their understanding and academic performance (Joshi et al. 2019; Bala et al. 2023). For this reason, recent research has expanded the scope of Learning Analytics (LA) into Multimodal Modal Learning Analytics (MMLA). Supported by advanced technologies such as deep learning, these approaches have achieved state-of-the-art results in predicting academic performance or knowledge. In particular, facial biometric signals have demonstrated a clear relationship with factors like attention, stress, and emotions, highlighting their potential for optimizing adaptive and personalized learning (Hernandez-Ortega et al. 2020a; Daza et al. 2023b; Kovanovic et al. 2023). VR offers a significant advantage in capturing facial expressions more easily and in an environment perceived as more realistic by the user, through VR headsets with facial sensors that allow detailed and continuous tracking of user interaction.
However, current public databases in e-learning have several limitations:
  • To our knowledge, there are no public databases specifically focused on VR environments for e-learning, as most are centered on traditional e-learning environments.
  • Most databases are not multimodal; they do not include biometric information or behavioral data.
  • The majority focus solely on learning assessment by collecting only responses to problems or questionnaires, without considering the continuous monitoring of the student’s Learning Session (LS).
These limitations hinder the study of AL systems in VR, particularly for data-driven multimodal models. They also prevent comprehensive comparisons of the advantages and drawbacks relative to traditional e-learning environments.
The contributions of this work are as follows: i) We present a survey of state-of-the-art AL and public e-learning databases; ii) We developed a conference visualization platform for VR, integrating an Auto QA system to assess student comprehension. This platform enables continuous monitoring by capturing metadata and facial expressions; iii) We designed an online session protocol, which includes an initial video lecture paired with a pretest, 10 video lectures with accompanying questions of varying difficulty levels, along with an automated note system using ChatGPT. This system allows students to quickly highlight information in VR environments using a color-coding scheme, facilitating review; iv) We have released an initial subset of the database, featuring 10 users in VR. Each LS lasts 2.5 hours, capturing metadata and facial features. The database was carefully designed to create a balanced dataset, specifically intended to train and evaluate data-driven interpretability models for AL; and v) We conducted preliminary experiments based on IRT systems to validate the dataset, using both local and global facial features collected during the lecture videos to predict whether the student has understood the content.
The remainder of the paper is organized as follows: Section 2 provides a summary of related work on AL with a focus on IRT and public databases in e-learning. Section 3 presents the VR conference visualization platform, the published database, and the proposed evaluation challenges for the community. Section 4 outlines the experiments and results. Finally, conclusions and directions for future work are discussed in Section 5.
2 Related Work
AL is an educational approach that personalises the learning experience for each student by dynamically adjusting content, pacing, and learning pathways based on individual needs and responses (Barbosa et al. 2024). This is relevant in e-learning, where students often receive identical resources without considering prior knowledge or skills, and direct interaction is limited (McCusker et al. 2013). AL has gained momentum in recent years, driven by advances in Artificial Intelligence (AI), Machine Learning, and data analytics.
2.1 Approaches in Adaptive Learning
Traditionally, learning style models, such as Felder-Silverman and Kolb (Felder 2002; Kolb et al. 2007), have been used to personalize teaching; however, their effectiveness has been questioned due to a lack of solid empirical evidence. These models classify students according to sensory and cognitive preferences (An and Carr 2017). In response, modern adaptive systems prioritize objective data analysis, such as performance analytics and biometrics.
Modern adaptive education systems aim to model an abstract representation of a student’s skill or knowledge level based on their performance and interactions with the educational system. This is a continuous process, as students evolve over time (Desmarais and Baker 2012). The technological foundation lies in LA, an area gathering and analysing data on student interactions and engagement, providing real-time insights for AL adjustments (Chatti et al. 2012). The field is advancing towards MMLA, integrating diverse data sources like biometrics, behaviors, and metadata to offer a more holistic view of the educational process (Hernandez-Ortega et al. 2020a; Daza et al. 2023b). Three student models have been widely developed:
  • Item Response Theory (IRT): A statistical model designed to evaluate both student ability ($\theta$) and question difficulty ($b$) based on their responses, making it highly suitable for creating adaptive tests. IRT models utilize estimation methods such as Maximum Likelihood or Bayesian approaches, often employing iterative techniques for parameter estimation. IRT models vary based on the parameters they incorporate. The most common version is the 1PL (Rasch model), which includes only the difficulty parameter ($b$) for each item.
    Since traditional IRT models typically assume that examinees’ abilities are randomly sampled from a standard distribution, recent research has explored combining IRT with deep learning techniques. A notable example is the Deep-IRT model (Tsutsumi, Kinoshita, and Ueno 2021), which introduces two twin neural networks: one for modeling student ability and another for item difficulty. The network inputs are one-hot vectors for the question and the student. This model outperforms traditional ones across diverse datasets, predicting unknown item responses more accurately from previous histories and proving effective even for small datasets. However, it may encounter challenges with sparse data.
    Recent studies have adapted the 1PL model with deep learning to incorporate biometric signals, such as facial expressions, for predicting an additional parameter: the student state. One prominent example is the SAD-IRT model (Zhou, Suzuki, and Kumano 2024), which adjusts item difficulty dynamically based on students’ cognitive and emotional states. Built upon Deep-IRT, SAD-IRT adds a branch with spatial (MARLIN) (Cai et al. 2023) and temporal feature (Temporal Convolutional Network, TCN) extractors. This branch processes facial images captured while students solve test items to derive the student state parameter. Evaluated on a dataset of 20 participants and 1,000 responses, SAD-IRT outperformed both the DEEP-IRT model and facial video-only models.
    Dataset Users VR Bio. LS Description
    ASSISTments 46,674 No No Eval. only Math responses (1 year); thousands of questions
    Junyi 247,606 No No Eval. only Math responses; 722 questions
    EdNet (KT4) 297,915 No No Logs & Eval. TOEIC exam responses with logs of consulted material (15 months); thousands of questions
    SAD-IRT 20 No Yes Eval. only 17 hours of assessments with facial tracking; 1,000 multiple-choice responses
    edBB 60 No Yes Eval. only 10 hrs assessment, focused on behavioral detection
    Ours 10 Yes Yes Class & Eval. 25 hours in VR, covering learning and assessment with facial tracking; 450 multiple-choice responses
    Table 1: Summary of public datasets in education (ASSISTments Foundation 2013; Chang et al. 2015; Choi et al. 2020; Zhou, Suzuki, and Kumano 2024; Hernandez-Ortega et al. 2020a), and SMARTe-VR. Bio. $=$ Biometrics, Eval. $=$ Evaluation, LS $=$ Learning Session and Logs $=$ Recorded material consulted by students.
  • Knowledge Tracing (KT): Models the evolution of a student’s knowledge by updating an internal ’knowledge state’ based on responses over a sequence of questions, using techniques like the Hidden Markov Model. Unlike IRT, KT focuses on sequential learning rather than static skills, without accounting for question difficulty or individual ability. Advanced KT models, such as Deep KT, leverage deep learning for enhanced response prediction accuracy, though they lack the interpretability of IRT (Zhou, Suzuki, and Kumano 2024). Combined IRT-KT methods have shown promising results, alongside recent efforts to integrate biometric and behavioral data to capture affective states (Ghosh, Heffernan, and Lan 2020).
  • Biometric and Behaviour-Based Learning Recognition (BBLR): These emerging models assess the student’s level of understanding during the session, without requiring constant feedback through answers. They integrate biometric and behavioral data—such as facial expressions (Zhou, Suzuki, and Kumano 2024), eye blinks (Daza et al. 2020, 2024c), keystroke dynamics (Morales et al. 2015, 2016), heart rate (Hernandez-Ortega et al. 2020b), head pose (Daza et al. 2024a; Becerra et al. 2024) and more—to estimate cognitive load (Daza et al. 2022, 2023a, 2024b), emotions (Bala et al. 2023), etc., linking these states to learning levels. Unlike traditional approaches such as IRT and KT, these models enable real-time feedback, adjusting content and pacing according to the student’s state. Their main challenge lies in a lack of interpretability due to complex deep learning architectures, which do not always clearly interpret the factors contributing to academic performance.
2.2 Adaptive Learning Platforms
Several educational platforms have adopted AL models. For example, Duolingo uses IRT models to optimize language learning (Maris 2020), and platforms like ASSISTments and Coursera employ KT models or similar methods (Feng, Heffernan, and Koedinger 2009). Additionally, public institutions have developed their own AL systems, including:
Learning Evidence Analytics Framework (LEAF): Developed at Kyoto University (Ogata et al. 2022), LEAF applies a learning analytics approach to collect, analyse, and visualise educational data. It includes tools such as BookRoll, Log Pallet, and LAView for real-time data visualization and feedback. Through AI-driven big data analysis, LEAF provides personalised learning support, enabling immediate feedback for students and instructors.
edBB: Developed at the Autonomous University of Madrid (Hernandez-Ortega et al. 2020a; Daza et al. 2023b), edBB focuses on learning recognition through biometric and behavioral analysis, using CNN and RNN models to estimate attention, heart rate, and more. Integrated with M2LADS (Becerra et al. 2023), it offers instructors a comprehensive view of student progress.
2.3 Public State of the Art Database
Public datasets in education mainly focus on student responses and assessment (see Table 1), with limited incorporation of biometric data and without VR. For example, the ASSISTments (ASSISTments Foundation 2013) and Junyi (Chang et al. 2015) datasets provide responses to math questions. Similarly, EdNet dataset (Choi et al. 2020) includes question responses and incorporates records of materials consulted by students during TOEIC exam preparation. All of these datasets lack both biometric and VR monitoring.
Few datasets explore multimodality with biometrics. The SAD-IRT dataset (Zhou, Suzuki, and Kumano 2024) includes facial tracking during literacy tasks, while edBB (Hernandez-Ortega et al. 2020a) records biometric signals such as heart rate, EEG and facial video for behavior analysis. However, these datasets focus on PC-based assessment.
In contrast, SMARTe-VR uniquely combines VR, learning, and assessment, with biometric monitoring, making it a valuable resource for AL system development.
3 SMARTe-VR: Platform for Adaptive E-Learning in Virtual Reality
SMARTe-VR platform is designed for e-learning sessions through VR, enabling instructors to upload their educational content while students experience it in an immersive environment that simulates a lecture hall with a screen, similar to works such as (Nagao 2023). The platform provides a VR environment that facilitates student interaction with the system, monitors learning data, and captures facial expressions in real-time. This feature allows detailed analysis of the student’s emotional and cognitive state. Additionally, it integrates learning assessment tools, enabling instructors to adjust content based on student performance.
3.1 Platform Features
VR Interaction: The immersive environment simulates a virtual lecture hall, allowing students to zoom in/out by moving their head, adjust brightness, and interact using VR controllers or a mouse. Features include answering ques- tions, opening textbook, highlighting text, and controlling video playback (pause, fast-forward, etc.).
Function Color
Marks learned content. Yellow
Highlights content for review. Blue
Indicates content not understood. Red
Marks key content for memorization. Green
Table 2: Functions for tagging and highlighting with their associated Colors in SMARTe-VR
Structure and Assessment Tool: Instructors can structure LSs using lecture videos of various lengths and content. It also features an automatic assessment system (Auto QA) that can be activated at different stages. Before the session: Preliminary questions assess initial student knowledge. After each video: Comprehension of the video content is assessed.
This method, used by platforms like Coursera and edX, is supported by research showing that a few well-crafted questions effectively measure real-time understanding (Cummins, Beresford, and Rice 2015). These assessments support adaptive models such as IRT and KT, enabling the platform to be used for intelligent AL.
Student Feedback: After each Auto QA intervention, students receive their results and a comparison with the top-performing student. At the session’s end, a score summary is provided. Continuous feedback and performance comparisons motivate students, helping them track progress, identify focus areas, and optimize learning using notes and resources (Van der Kleij, Feskens, and Eggen 2015).
Textbook Interaction System: SMARTe-VR generates summaries of main ideas from each lecture video using subtitles and a Large Language Model (ChatGPT). Instructors can review and edit these summaries to create a textbook. Students can view and highlight the text anytime, even during Auto QA questioning. To encourage interaction and collect learning data, two methods are used:
  • Textbook Highlighting: Students can mark text with a color-coded system (see Table 2) indicating importance and comprehension level.
  • Quick Tagging Buttons: Four colour-coded buttons (see Table 2) enable students to tag video sections quickly. The platform labels these sections, providing valuable insights for instructors and learning analysis.
Studies show that visual interaction mechanisms improve content retention and comprehension. Highlighting key information is associated with higher retention and more efficient review (Ogata et al. 2022).
Device Features Information
VIVE XR Elite 90Hz
1920x1920 resolution
4 tracking cameras
speakers
Visual display and audio, spatial tracking, and motion detection, responsible for transmitting data to the PC.
VIVE XR Facial Tracker 30Hz
Dual-tracking eye camera
Mono-tracking face camera
Attachable to the VR headset, tracks facial expressions and movements, and detects interpupillary distance.
PC Mouse used as a controller
Data processing
Records: 51 facial features, timestamps, interaction buttons, video feedback tags, textbook interactions (e.g., consulted pages), lecture videos, questions, responses, selection history, scores, understanding labels, pretest results, content/question difficulty, and question concept tags.
Table 3: Overview of devices and data collected by the monitoring system.
3.2 SMARTe-VR Database [^1] Access: https://github.com/BLANCELIN1/SMARTe-VR-DB
We monitored LSs focused on the TOEIC, specifically Section 5, which assesses grammar comprehension. The TOEIC is a recognized English test used by companies to evaluate non-native speakers’ language skills in professional contexts. This study involved 10 Japanese students from Hiroshima University. The dataset comprises 450 answered questions and approximately 25 hours of learning data, including facial movements and expressions.
Ethical Considerations.
The database was collected in accordance with the Declaration of Helsinki and approved by the Ethics Committee. Learners were briefed on the procedure and data acquisition, providing informed consent. Due to the extended monitoring, participants received an incentive, with an additional reward for top performers.
Devices.
Table 3 lists the devices and data types captured. The setup included the following components (see Fig. 1):
VR Headset: We used the VIVE XR Elite, designed for educational and business applications, along with the VIVE XR Facial Tracker. This accessory uses infrared cameras to measure interpupillary distance for optimal viewing and enable facial tracking. Using the OpenXR SDK, we monitored up to 51 facial features at 30 Hz, ensuring system stability and preventing overheating. Captured facial features include movements and expressions, quantified from 0 to 1 in intensity. These features include eye movements and blinking, jaw opening, mouth and lip expressions (e.g., smiles) cheek gestures (e.g., puffing), and tongue movements.
Personal Computer: It connected to the VIVE XR Elite manages the virtual environment, collects and analyzes sensor data, and captures learning records, ensuring optimal performance without headset overheating. Although VIVE XR Elite controllers are supported, we chose the mouse as the VR controller because: 1) Students remained seated, simulating a classroom, making it more practical; 2) Test feedback indicated it was more comfortable for interacting with the SMARTe-VR interface and highlighting text; 3) This approach allows for better comparisons to traditional e-learning setups for future research.
Protocol and Tasks
The goal was to design a monitoring system for real VR LSs, covering both learning and assessment processes while tracking learning metadata and facial expressions. This protocol aims to aid the research community in developing frameworks like IRT, KT, MMLA, and BBLR, filling the gap in databases with these features.
The content of each lecture video and its questions was designed to create a balanced dataset, trying an equal number of videos were labeled as understood or not understood. Ten 8-minute TOEIC exam lecture videos, explained in Japanese, were selected from YouTube and varied in difficulty according to TOEIC guidelines. Each video had three multiple-choice questions with four options, focused on sentence completion to assess grammar and vocabulary. A video was labeled understood if a student answered at least 2 out of 3 questions correctly. The questions were from a TOEIC dataset on the ProProfs platform, which includes data on response accuracy and participant numbers.
Using the sample size estimation formula, we determined the minimum number of responses needed to represent 99% of the population with a 1% margin of error, enabling the classification of questions by difficulty and content. Questions were categorized as easy (80% correct response rate), medium (50%), and difficult (20%). The question structure per video was: easy-level videos had 2 easy and 1 medium question; medium-level had 3 medium questions; difficult-level had 2 difficult and 1 medium question.
To verify difficulty accuracy, a preliminary test with similar students was conducted, collecting subjective evaluations on each session’s difficulty and its questions.
Each student participated in a 120-150 minute LS consisting of 11 lecture videos and 45 questions, similar to a university class. The sessions included:
Before Starting the Session:
Calibration and Student Information: The supervisor explained the session protocol and verified data capture. A 3D facial model was monitored throughout to confirm recording of facial movements and expressions. Privacy Policy: The students reviewed a document detailing the data collection and its purpose, with the option to accept or decline participation. Video Lecture 0 and Auto QA Trial Phase: The students began with an initial lecture video that explained how to use the interface. At the end of this video, the Auto QA system was activated, presenting 5 questions without a time limit to help students become familiar with the interface. Auto QA Pretest: A pretest of 10 questions (3 easy, 4 medium, and 3 difficult) related to the upcoming lecture content was designed with two main objectives: i) to evaluate the students’ current knowledge level and ii) to serve as a baseline for IRT or similar systems.
Learning Session:
The session consisted of 10 lecture videos, each followed by 3 related questions. The structure was: i) Students watched an 8-minute video, interacting with the textbook and color-coded buttons; ii) The Auto QA was activated post-video, presenting 3 questions with a one-minute response time each, allowing students to review the textbook. This immediate questioning reduced potential memory issues; iii) Students received feedback and compared their performance to the top student; iv) Students could take a break if needed—most did not, but those who did paused for approximately 5 minutes; v) A 5-10 minute VR headset battery change occurred after the 5th video.
End of the Session:
The students received a total score and a comparison with the top-performing student. They provided feedback to the supervisor about their SMARTe-VR experience, which highlighted the following: i) Prolonged VR use did not cause motion sickness as participants remained seated in the VR; ii) Students reported restlessness over time due to limited interaction options; and iii) Some students reported discomfort when using the mouse in VR, as they were unable to see their hands.We conclude that integrating Mixed Reality (MR) and shortening sessions could enhance interaction, device usability (e.g., mouse, keyboard), and overall user experience.
Challenges
Figure 1: Example setup and data captured by the SMARTe-VR platform, including the proposed understanding detection models (SMART-TCN and SMART-MLP) for database validation.
We propose 5 challenges for the research community, each with specific target and input data. The objective is to leverage the SMARTe-VR database to develop AL models that estimate student performance, knowledge, or behavior. The proposed challenges include:
Challenge 1 - Understanding Estimation: Predicting student comprehension of educational content is valuable for AL. We propose estimating if a student understood lecture content using binary labels (understood, not understood). Target: Binary classification to predict student understanding in VR based on multimodal data. Inputs: Facial features, metadata, pretest results, and content difficulty.
Challenge 2 - Response Prediction: Predicting student responses using IRT or KT models with multimodal data. Proposed scenarios: i) Response data only, ii) Adding multimodal data captured during the response (one minute/question), and iii) All data from content and responses. Target: Predict correct answers. Inputs: 450 questions, facial features, metadata (e.g., response time), question concept tags, content and question difficulty, and pretest results.
Challenge 3 - Student Feedback Prediction: The ability to anticipate how a student perceives content (e.g., important, learned/not learned, to memorize) is crucial for AL. This challenge seeks to create a model that uses color-coded button interactions from VR sessions as ground truth to predict feedback. Target: Predict student feedback on their understanding and the perceived relevance of the content. Inputs: Facial features, learning metadata, pretest results, and lecture videos.
Challenge 4 - Response Time Detection: Predicting if a student will respond before the time limit can help gauge confidence, detect fatigue or inattention, and adjust workload or response time dynamically. Only biometric data and metadata from the first 20 seconds are permitted. Target: Predict if a student will respond on time. Inputs: Facial features, metadata (e.g., selection path history), content and question difficulty, and pretest results.
Challenge 5 - Performance Prediction: Student performance was measured by accuracy (percentage of correct responses), enabling classification into three levels (low, medium, high). We propose estimating final performance across three levels using biometric data and session metadata without incorporating response data. Indicators such as attention, emotional state, and fatigue from facial features, along with learning metadata, can help in predicting performance. Target: Performance. Inputs: Facial features, learning metadata, content difficulty, and pretest results.
4 Experiments and Results
4.1 IRT Models
We tested two IRT (1PL) model architectures based on (Zhou, Suzuki, and Kumano 2024) to validate the dataset, aiming to predict student understanding during video lectures (Challenge 1). These models used facial features within a time window ${W_{l}\in\{1,\ldots,8\}}$ minutes, along with one-hot vectors for the user and the questions. The binary understand label served as the ground truth for prediction. Fig. 1 shows a schematic of these models:
SMART-TCN model: Based on Deep-IRT (see Section 3), we used two twin Multilayer Perceptrons (MLPs) to estimate question difficulty ($b$) and student ability ($\theta$) separately. One MLP processed a one-hot vector representing the question to obtain $b$, while the other processed a one-hot vector identifying the student to estimate $\theta$. The student state ($\phi$) was based on SAD-IRT, using a TCN that processed facial features extracted from video lectures via the VR headset. The TCN was configured with 4 layers (32 to 64 neurons), a filter size of 3, and an expansion rate of 1, 1, 2, and 4, optimized to capture subtle changes in expressions held for extended periods. The TCN output was then connected to a fully connected layer with an output layer using a Tanh function to obtain the state parameter. Finally, the system estimated the probability of understanding as $P(\text{Understand})=\frac{1}{1+e^{-(\theta-(b+\phi))}}$.
SMART-MLP Model: This model was similar to the SMART-TCN but replaced the TCN with three parallel MLPs to estimate each parameter independently. Each MLP had two hidden layers, the first with 128 neurons and the second with 64 neurons, both using ReLU activation, followed by an output layer with a Tanh function. This model used the global facial features as input to estimate $\phi$.
Both models were trained using the Adam optimizer with a learning rate of 1e-03, a batch size of 64, and a total of 200 epochs. A cross-entropy loss function was applied for binary prediction of understanding.
Local Features
For SMART-TCN model, we used local facial features, with the VR headset providing 30 sets of 51 features per second. The resulting vector is $\textbf{f}_{\mathrm{L}}\in\mathbb{R}^{30\times 51\times W_{l}}$, dependent on the selected time window $W_{l}$.
Global Features
For SMART-MLP model, we used global facial features, specifically 8 global features for each facial feature. These are the maximum, minimum, mean, standard deviation, median, kurtosis, skewness, and rate of change. The resulting vector is $\textbf{f}_{\mathrm{G}}\in\mathbb{R}^{8\times 51}$.
4.2 Experimental Protocol
SMARTe-VR is a nearly balanced dataset regarding conference understand labels, with 55% of sessions marked as understood. Specifically, 83% of easy, 45% of medium, and 40% of difficult sessions are marked as understood. Each user has 10 labels for each learning session, totaling 100 labels across all users.
We used a “leave-one-out” cross-validation approach, where each user is left out once for testing while the remaining users are used for training; this process repeats for each user. However, even with leave-one-out, the model is always trained on video lecture 0 and the 10 pretest questions, including those from the left-out user, to estimate their initial ability values. The decision threshold is set at the point where the False Positive and False Negative rates are equal.
Figure 2: Comparison of accuracy across different time windows ($W_{l}$) for our proposed models and the Traditional IRT model on the SMARTe-VR database.
4.3 Results
Figure 2 presents the results obtained for the proposed models compared with traditional IRT across various time windows ${W_{l}\in\{1,\ldots,8\}}$ minutes.
Results indicated that with short time windows, traditional IRT outperformed the two deep learning-based models, which struggled in this range. However, from 4 minutes onward, facial features began to provide valuable insights, enhancing both models’ ability to predict student state and understanding beyond the capabilities of traditional IRT. Both models showed improved performance as $W_{l}$ increased, suggesting that facial features more accurately represented the student’s state over longer periods. With time windows of several minutes, the models proposed were better able to capture behavioral patterns in facial expressions. Additionally, given that Japanese students tend to maintain more neutral and consistent facial expressions compared to other cultures, facial changes were likely subtler and less frequent. This suggests that, with shorter time windows, the collected facial signals may not have sufficiently reflected the student’s comprehension state. SMART-MLP, based on global features, achieved the highest accuracy of 85% in the 8-minute window. Both models demonstrated similar performance across all time windows, though SMART-MLP performed slightly worse in shorter windows, likely due to the limitations of global features in capturing rapid patterns.
Content Difficulty
Model Easy Medium Hard
SMART-MLP ($W_{l}=8$) 0.828 0.868 0.851
Traditional IRT 0.682 0.741 0.702
Table 4: Accuracy obtained for the SMART-MLP model with a $W_{l}=8$ minutes and for the traditional IRT model, evaluated across different content difficulty levels labeled in the SMARTe-VR dataset.
Finally, for the SMART-MLP model in the 8-minute time window, which achieved the best results, we analyzed its accuracy by the content difficulty levels provided in the dataset (see Table 4). The SMART-MLP model outperformed the traditional IRT model across all content. Both models showed a similar trend, being more effective in detecting understanding for medium-level content, followed by difficult-level content, with the lowest accuracy for easy-level content. With more medium-difficulty sessions in the dataset, SMART-MLP can more easily identify patterns. Additionally, IRT models benefit from medium-difficulty items, as the variability in responses allows the model to better discriminate between skill levels, crucial for accurate estimates.
5 Conclusion and Future Work
In this work, we have: i) reviewed adaptive learning models and datasets; ii) introduced a VR platform for immersive conference visualization in e-learning, specifically designed for adaptive learning, which monitors both biometric data and learning metadata; iii) released the 5 public VR-based educational dataset; iv) proposed five research challenges tied to this dataset; and v) designed two architectures for understanding detection to support dataset validation.
We believe VR will transform education. SMARTe-VR provides a comprehensive VR solution equipped with tools such as Auto QA, textbook creation, and biometric monitoring, enhancing educator capabilities and enabling AI-driven advancements. The SMARTe-VR dataset and its associated challenges offer a valuable resource for advancing AL models in VR, supporting research in IRT, KT, and MMLA
SMART models offer a novel approach based on state-of-the-art IRT models, with a focus on understanding detection during learning. By leveraging facial biometrics, they enhance adaptability, enable real-time adjustments based on student states, and reduce reliance on response data. These methods validated our dataset with 85% accuracy, although further research is needed to ensure broader applicability.
In future work, we plan to expand the dataset with more users, additional questions per video, a broader range of lecture videos, and further sensor integrations to create a more robust resource. We also plan to monitor students in e-learning settings for direct comparisons.
Furthermore, MR will be integrated into the platform to enhance the user experience. To address the challenges, we will share models with the community and refine SMART-MLP and SMART-TCN models to detect understanding and include attention and fatigue detection. Ultimately, our goal is to incorporate these models into SMARTe-VR, creating a comprehensive adaptive learning experience that leverages the potential of VR-based education.
6 Acknowledgments
This study has been supported by project BBforTAI (PID2021-127641OB-I00 MICINN/FEDER). A. Morales is also supported by the Madrid Government under the Multiannual Agreement with UAM in the line of Excellence for the University Teaching Staff (V PRICIT). This study is particularly grateful to the experimenters from Nagoya University and Hiroshima City University for their significant contributions to data collection and experiment implementation. Additionally, we extend our thanks to both universities for providing experimental equipment and technical support, which were indispensable for the successful completion of this research. Without their support, this study would not have been possible.